# ElasticTok: Adaptive Tokenization for Image and Video

### v1

@@ 2.2 Inference
  Required: Input video $x$. Tokens per block $B$.
@@
     Initialize masks $m_{i}\leftarrow\mathbf{0}_{B}$
@@
  Required: Input sequence $x$. Tokens per block $B$.
@@
     $\ell_{i},m_{i},c\leftarrow\text{SearchAlgo}(x_{iB:(i+1)B},c,t,D,E)$
     Encode $z_{i}\leftarrow E(x_{iB:(i+1)B},m_{i},c)\odot m_{i}$
@@ 3.1 Experimental Setup
Model details. We train 200M parameter discrete and continuous autoencoders on images and video sequences. For learning discrete representations, we use FSQ (Mentzer et al., 2023), while for continuous representations, we use a VAE (Kingma, 2013). Both models employ ViT encoders and decoders. Similar to prior works (Esser et al., 2021), we optimize both a pixel-wise reconstruction loss (MSE) and a perceptual loss (LPIPS) (Johnson et al., 2016). Notably, we do not use a GAN (Goodfellow et al., 2014), as we find it more difficult to stably train when progressively extending to much longer sequences (e.g. 1000+ frames), which we leave to future work, as it is not the main focus of this paper.
@@
Figure 3: ElasticTok adaptively encodes image and video to variable length outputs based on the complexity of the input data. (Top) Ground-truth video frames. (Middle) Reconstructed frames with varying token usage. (Bottom) The bottom section depicts how ElasticTok dynamically adjusts token allocation over time, with the percentage of tokens used correlating to different content complexities in the video.
@@ 3.2 Main Results
Figure 4 shows quantitative comparisons between our method and different fixed token baselines trained at each token length. In order to show the benefits of elastic token representations, we compute a quantitative metric that measures the percentage of images or videos blocks in which a model satisfies a specified reconstruction (MSE) threshold. As shown, our method can leverage its elastic representations to achieve similar reconstruction satisfactory percentages compared to baselines while using 1.3x - 5x fewer tokens, depending on the given threshold. Generally, more lax (0.015) reconstruction thresholds present larger performance improvements due to more room to use fewer than max tokens. In contrast, more strict (0.001) thresholds usually almost always require all tokens. Different thresholds may be useful for different cases, where more lossy encodings can be used for simple VQA tasks, while more accurate reconstructions are useful for tasks such as image / video generation. Table 3.2 shows similar results for the continuous variant of our model, compared against a single baseline model fixed to $50\%$ token usage.
Table 3.2 shows results comparing ElasticTok against a quality baseline. In which we first run inference using our model with a given threshold, and then compute which token length baseline is needed to match the quality (worst-case MSE over video blocks) of our model. Note that our model can use a different number of tokens each video block, where as the baseline uses the same number of tokens for all blocks. The adaptivity of our model helps match baseline model quality with lower average token usage.
@@
Takeaway: ElasticTok offers significant advantages in image and video reconstruction by using variable-length codes that depend on content complexity. Qualitatively, simpler images require fewer tokens, while more complex ones need more tokens for accurate encoding. Quantitatively, elastic tokenization achieves similar reconstruction quality compared to fixed token baselines while using up to 5x fewer tokens
Figure 4: Performance comparison between baseline and ElasticTok on ImageNet and Video. The y-axis shows the percentage of samples that satisfy the reconstruction threshold, while the x-axis represents the percentage of tokens used. (Left) On image, ElasticTok achieves a 3.5x and 1.3x efficiency boost at different reconstruction thresholds. (Right) On video, ElasticTok shows a 5x and 2.4x improvement over the baseline, maintaining superior performance while using fewer tokens. Figure 10 in Appendix D shows reference examples of reconstruction quality for an image at different thresholds.
Table 1: Performance comparison between baseline and ElasticTok across different reconstruction thresholds with the continuous token model. Percentage shows the proportion of video blocks that satisify the given reconstruction threshold. Our model can satisfy more video blocks.
{NiceTabular}
@lccc@ Threshold
$\mathbf{0.001}$ $\mathbf{0.003}$ $\mathbf{0.015}$
Baseline $40\%$ $78\%$ $99\%$
Ours $\mathbf{50\%}$ $\mathbf{88\%}$ $\mathbf{99\%}$
Table 2: Comparing performance of our Quality Baseline and ElasticTok at different reconstruction thresholds. Our model can use fewer tokens to reach the same quality (match worse-case MSE) as the baseline model.
{NiceTabular}
@llll@ Threshold
0.001 $\mathbf{0.003}$ $\mathbf{0.015}$
Baseline $78\%$ $57\%$ $15\%$
Ours $\mathbf{75\%}$ $\mathbf{37\%}$ $\mathbf{9.4\%}$
Figure 5: Loss progressively declines as more tokens are used. The top row illustrates the impact on text clarity, while the bottom row shows the effect on image sharpness. The graphs on the right quantify the reconstruction loss relative to token usage percentage, showing a rapid decline as more tokens are consumed.
Figure 6: Progressive performance increase with more frames. Performance improves with increasing sequence length, peaking around 100 frames before a slight decline. (Note the log scale for the x-axis)
@@ 3.3 Downstream Tasks
Figure 7: The accuracy and compute trade-off with varying percentages of tokens used. This allows users to adjust the accuracy based on computational budget.
@@
Table 3: Comparison of our method with baseline on image and video benchmarks. Our method can match the performance of the baseline trained on a fixed number ($100\%$) of tokens. However, baseline models are restricted to a fixed token output, and require full pretraining a new model for every possible token length, whereas ElasticTok only requires a single model to generalize to all token lengths.
@@ 3.4 Inference
Refer to caption Figure 8: Comparison of different loss functions for inference. Inference with a CLIP model (cosine distance) prioritizes textual reconstruction (bottom) while deprioritizing other detailed visual features (top).
Table 4: Comparison of inference methods showing their respective error rates, number of function evaluations (NFEs). Note that while Full and Binned Search are more computationally expensive, they could also benefit more from parallel function evaluations if compute is available.
3.5 Frequency Analysis
@@
Figure 9: Comparison of token usage versus frequency magnitude in single-block and multi-block frequency analysis. Both scatter plots show a strong positive correlation between frequency magnitude and token usage in a single-block setting a multi-block setting. The red lines represent the linear regression fits for each case.
@@ 4 Related Work
Adaptive Representation. There have been a lot of research studies on learning adaptive or ordered representations. Early work on nested dropout (Rippel et al., 2014) proposes using dropout (Srivastava et al., 2014) in the context of autoencoders to learn ordered representations. In this approach, an index is sampled from a prior distribution, and all units with an index larger than the sampled one are dropped. Similar to our ElasticTok, this method imposes an inherent ordering on the representation dimensions. Units that are dropped less frequently encode more important information, while those that are dropped more often encode less critical details. Other works study adaptive architectures based on context size (Kim and Cho, 2020), slimmable neural networks that train a network by sampling multiple sub-networks of different channel numbers simultaneously, where the weights are shared among different widths (Lee and Shin, 2018, Yu et al., 2018), adaptive width and depth in BERT (Hou et al., 2020), or dropping random layers during training to increase robustness to pruning (Fan et al., 2019, Huang et al., 2016). Dieleman et al. (2021) learns a variable-rate representation on audio applied to speech. Transframer (Nash et al., 2022) uses variable-length DCT representations with transformers for general image and frame-level video prediction tasks. Matryoshka representation learning (Kusupati et al., 2022) explores adding nested substructures inside the standard Transformer block. However, none of the previous works consider learning elastic representations in autoencoders for long sequences like videos. Our work provides a scalable solution for representing videos with elastic representations.
Tokenization. Representing visual images with a set of discrete tokens has been extensively studied, such as in VQVAE (Van Den Oord et al., 2017, Razavi et al., 2019) and VQGAN (Yu et al., 2021). Recent research has highlighted several shortcomings of traditional tokenization methods, including low vocabulary utilization. In response, alternative approaches like FSQ have been explored (Mentzer et al., 2024, Yu et al., 2023b). These discrete visual tokens facilitate integration with next-token prediction in language models and multimodal models (Yu et al., 2023b, Liu et al., 2024a, Xie et al., 2024, Wang et al., 2024). Parallel to this, other studies have investigated the use of continuous embeddings for representing images and videos (Lin et al., 2023, Liu et al., 2023, Zhu et al., 2023). Other research proposes next-scale prediction (Tian et al., 2024), which employs hierarchical multi-scale modeling to first capture coarse details, followed by finer ones. However, this approach requires a manually defined hierarchy and uses a fixed number of tokens. Our work, while complementary to these efforts, focuses on adaptive representation and can be directly applied to both discrete and continuous approaches.
Compression. Adaptive learning for images and video have also been well-studied in the context of variable-rate compression. JPEG (Christopoulos et al., 2000) remains one of the most popular lossy image compression algorithms in the world, using a combination of DCT with quantization followed by entropy coding. For video, codecs such as H264 (Wiegand et al., 2003), H265 (Sze et al., 2014), VP9 (Mukherjee et al., 2015), and AV1 (Han et al., 2021) are popularly used compression algorithms. Prior works have additionally explored extending neural networks to learn more effective compressors. Theis et al. (2022) and Minnen et al. (2018) introduce methods that leverage CNNs to learn effective variable bit-rate compression algorithms over images competitive with JPEG. Li et al. (2023a), Mentzer et al. (2022), and Li et al. (2021) extend neural networks to learn how to effectively compress videos. While similar to our work in that these methods also learn adaptive encodings, these compression methods are more difficult to utilize for downstream training and generation tasks due to lack of direct compatibility. In contrast, our work builds upon existing tokenization strategies (VAE, FSQ) that have shown to work well for such downstream tasks, as well as take advantage of the adaptive representations to better utilize compute.
@@ 5 Discussion and Conclusion
In this work, we propose an elastic representation approach address the inefficiencies of traditional video encoding approaches by introducing an adaptive encoding method that selectively encodes new information based on the context of previous frames. By dynamically allocating resources, it reduces computational costs while maintaining high-quality video representation. The proposed technique of dropping tokens at the end of each sequence allows the model to prioritize essential information, ensuring scalability and efficiency during inference. Our ElasticTok method demonstrates strong performance in both video representation and downstream tasks.
Lastly, we identify some limitations of our method, as well as possible directions for future work.
  • Masking schemes: Our model currently slightly underperforms baselines at the tail ends of encoding lengths, which we hypothesize may be partially due to conflicting representations that the model needs to learn in each case (low frequency, global encodings vs high frequency, local encodings). Preliminary investigations showed that changing the way tokens are distributed (as opposed to our method of left-aligned tokens) improved reconstruction accuracy. Future work may also explore learnable encoding masks to dynamically select which tokens to keep for each input.
  • Other temporal modalities: Although our primary focus was on image and video, our method is generally modality-agnostic and can be extended to any temporal data, such as audio and decision-making trajectories (e.g., state, action, reward).
  • Conditional encoding and different training objectives: Our focus was on leveraging temporal encoding to achieve more efficient tokenization while retaining high reconstructed visual quality. However, our method can be explored more broadly as a means of encouraging meaningful lossy encodings. For example, robotics models can learn text-conditioned visual encoders that capture only task-relevant information from an image. In other scenarios, different reconstruction objectives could prioritize facial or textual reconstruction while ignoring background reconstruction accuracy.
@@ Acknowledgments
We gratefully acknowledge the Google TPU Research Cloud (TRC) program for providing access to TPUs. We extend our thanks to Xinyang Geng, Jimmy Shi, Tim Rocktäschel, Joao Carreira, and Satinder Singh for their valuable discussions and insightful feedback on an earlier version of this paper, and to Doina Precup for valuable support of this work. Pieter Abbeel holds concurrent appointments as a Professor at UC Berkeley and as an Amazon Scholar. This paper describes work performed at UC Berkeley and is not associated with Amazon.
@@ Appendix A Model configuration
ElasticTok-VAE (Continuous) ElasticTok-FSQ (Discrete)
Parameters 210M 210M
Frame Resolution $256\times 256$ $256\times 256$
Block Size ($M_{max}$) 2048 tokens (4 frames) 4096 tokens (4 frames)
$M_{min}$ 128 tokens 256 tokens
Max Number of Frames 1024 1024
Patch Size (T, H, W) (2, 8, 8) (1, 8, 8)
Hidden Size 1024 1024
FFN Size 2048 2048
Encoder Layers 10 10
Decoder Layers 10 10
RoPE Theta 5000000 50000000
Max Sequence Length 512K 1M
FSQ Dims N/A 8, 8, 8, 5, 5, 5 (64k codes)
VAE Dim 8 N/A
KL Weight 1e-8 N/A

### v2

@@ 1 Introduction
2 Background
2.1 Blockwise RingAttention
Blockwise RingAttention (Liu et al., 2024b) is a training technique for efficiently scaling transformers on long-context data. Each rank along the sequence parallel sharding axis retains non-overlapping, equal sizes slices of the full sequence. During an attention operation, each sequence parallel rank calculates its own corresponding $Q_{i},K_{i},V_{i}$, and computes its own portion of the full attention as: $\text{softmax}(Q_{i}^{T}[K_{1},\dots,K_{S}])[V_{1},\dots,V_{S}]$, where $S$ is size of the sequence parallel dimension. Instead of naively all-gathering $K,V$, RingAttention proposes to iteratively pass $K_{i},V_{i}$ blocks in a ring-structure, where rank $i$ will pass its current key-value block to rank $(i+1)\mod S$. Importantly, RingAttention enables overlap between compute and communication by computing the partial attention over the current block while communicating the next block. In this paper, we leverage RingAttention to scale our method to long sequences of videos.
@@ 3.2 Inference
Further ablations (Table F in Appendix F) suggest that we can additionally achieve 2x inference speedup at a slight cost to encoding quality by not conditioning the mask on the encoder – this would require only running the encoder once, and then performing iterative search through masking the encoding and only running the decoder. This may be a better choice depending on the size of the downstream pretrained model.
@@
  Required: Video $x$. Patch size $T_{p}\times H_{p}\times W_{p}$.
  Required: Tokens per block $Z$
@@
  Required: Min / Max token lengths $M_{min},M_{max}$
  // batch size, timesteps, height, width, channels
  // $x$ is $B\times T\times H\times W\times C$
@@
     Initialize masks $m_{i}\leftarrow\mathbf{0}_{Z}$
@@
  Required: Video $x$. Patch size $T_{p}\times H_{p}\times W_{p}$.
  Required: Tokens per block $Z$.
@@
  // batch size, timesteps, height, width, channels
  // $x$ is $B\times T\times H\times W\times C$
@@
     $\ell_{i},m_{i},c\leftarrow\text{SearchAlgo}(x_{iZ:(i+1)Z},c,t,D,E)$
     Encode $z_{i}\leftarrow E(x_{iZ:(i+1)Z},m_{i},c)\odot m_{i}$
@@ 4.1 Experimental Setup
Model details. We train 200M parameter discrete and continuous autoencoders on images and video sequences. For learning discrete representations, we use FSQ (Mentzer et al., 2023), while for continuous representations, we use a VAE (Kingma, 2013). Both models employ ViT encoders and decoders, consisting of a single patchify strided convolutions (the transpose for the decoder), followed by transformer blocks identical to those in LLaMA 2 (Touvron et al., 2023). Attention is applied in a block-causal manner, with block size equal to the number of tokens per video block. The encoder is additionally conditioned on the token mask (represented as a binary mask) by replacing 0’s and 1’s with different learned embedding vectors, and then added to the video patchify output hidden states. The bottleneck representations are computed through a linear projection on the encoder output. Similar to prior works (Esser et al., 2021), we optimize both a pixel-wise reconstruction loss (MSE) and a perceptual loss (LPIPS) (Johnson et al., 2016). Notably, we do not use a GAN (Goodfellow et al., 2014), as we find it more difficult to stably train when progressively extending to much longer sequences (e.g. 1000+ frames), which we leave to future work, as it is not the main focus of this paper.
@@
Evaluation Search Algorithm Unless otherwise noted, all results use the Binary Search algorithm as described in Section 3.2.
Figure 3: ElasticTok adaptively encodes image and video to variable length outputs based on the complexity of the input data (using ElasticTok-VAE). The top rows shows examples of ElasticTok on images. Below shows a video example with: (Top) Ground-truth video frames. (Middle) Reconstructed frames with varying token usage. (Bottom) The bottom section depicts how ElasticTok dynamically adjusts token allocation over time, with the percentage of tokens used correlating to different content complexities in the video.
Figure 4: Performance comparison between baseline and ElasticTok -FSQ on ImageNet and Video. The y-axis shows the percentage of samples that satisfy the reconstruction threshold, while the x-axis represents the percentage of tokens used. (Left) On image, ElasticTok achieves a 3.5x and 1.3x efficiency boost at different reconstruction thresholds. (Right) On video, ElasticTok shows a 5x and 2.4x improvement over the baseline, maintaining superior performance while using fewer tokens. Figure 10 in Appendix D shows reference examples of reconstruction quality for an image at different thresholds.
@@ 4.2 Main Results
Figure 4 shows quantitative comparisons between our method and different fixed token baselines trained at each token length. In order to show the benefits of elastic token representations, we compute a quantitative metric that measures the percentage of images or videos blocks in which a model satisfies a specified reconstruction (MSE) threshold. As shown, our method can leverage its elastic representations to achieve similar reconstruction satisfactory percentages compared to baselines while using 1.3x - 5x fewer tokens, depending on the given threshold. Generally, more lax (0.015) reconstruction thresholds present larger performance improvements due to more room to use fewer than max tokens. In contrast, more strict (0.001) thresholds usually almost always require all tokens. Different thresholds may be useful for different cases, where more lossy encodings can be used for simple VQA tasks, while more accurate reconstructions are useful for tasks such as image / video generation. Table F shows similar results for the continuous variant of our model, compared against a single baseline model fixed to $50\%$ token usage.
@@
Figure 5: Loss progressively declines as more tokens are used (ElasticTok-FSQ). The top row illustrates the impact on text clarity, while the bottom row shows the effect on image sharpness. The graphs on the right quantify the reconstruction loss relative to token usage percentage, showing a rapid decline as more tokens are consumed.
Figure 6: Progressive performance increase with more frames (ElasticTok-FSQ). Performance improves with increasing sequence length, peaking around 100 frames before a slight decline. (Note the log scale for the x-axis)
Takeaway: ElasticTok offers significant advantages in image and video reconstruction by using variable-length codes that depend on content complexity. Qualitatively, simpler images require fewer tokens, while more complex ones need more tokens. Quantitatively, ElasticTok achieves similar reconstruction quality compared to fixed token baselines while using up to 5x fewer tokens
@@ 4.3 Downstream Tasks
Figure 7: The accuracy and compute trade-off with varying percentages of tokens used (ElasticTok-VAE). This allows users to adjust the accuracy based on computational budget.
@@
Table 1: Comparison of our method with baseline on image and video benchmarks (ElasticTok-VAE). Our method can match the performance of the baseline trained on a fixed number ($100\%$) of tokens. However, baseline models are restricted to a fixed token output, and require full pretraining a new model for every possible token length, whereas ElasticTok only requires a single model to generalize to all token lengths.
@@ 4.5 Frequency Analysis
Refer to caption Figure 8: Comparison of different loss functions for inference (ElasticTok-FSQ). Inference with a CLIP model (cosine distance) prioritizes textual reconstruction (bottom) while deprioritizing other detailed visual features (top).
Table 2: Comparison of inference methods showing their respective error rates, number of function evaluations (NFEs) (ElasticTok-FSQ. Note that while Full and Binned Search are more computationally expensive, they could also benefit more from parallel function evaluations if compute is available.
@@ 5 Related Work
Adaptive Representation. There have been a lot of research studies on learning adaptive or ordered representations. Early work on nested dropout (Rippel et al., 2014) proposes using dropout (Srivastava et al., 2014) in the context of autoencoders to learn ordered representations. In this approach, an index is sampled from a prior distribution, and all units with an index larger than the sampled one are dropped. Similar to our ElasticTok, this method imposes an inherent ordering on the representation dimensions. Units that are dropped less frequently encode more important information, while those that are dropped more often encode less critical details. Other works study adaptive architectures based on context size (Kim and Cho, 2020), slimmable neural networks that train a network by sampling multiple sub-networks of different channel numbers simultaneously, where the weights are shared among different widths (Lee and Shin, 2018, Yu et al., 2018), adaptive width and depth in BERT (Hou et al., 2020), or dropping random layers during training to increase robustness to pruning (Fan et al., 2019, Huang et al., 2016). Dieleman et al. (2021) learns a variable-rate representation on audio applied to speech. Transframer (Nash et al., 2022) uses variable-length DCT representations with transformers for general image and frame-level video prediction tasks. Matryoshka representation learning (Kusupati et al., 2022) explores adding nested substructures inside the standard Transformer block. Graves (2016) explores using recurrent neural networks to learn adaptive computation for learning tasks. However, none of the previous works consider learning elastic representations in autoencoders for long sequences like videos. Our work provides a scalable solution for representing videos with elastic representations. Tokenization. Representing visual images with a set of discrete tokens has been extensively studied, such as in VQVAE (Van Den Oord et al., 2017, Razavi et al., 2019) and VQGAN (Yu et al., 2021). Recent research has highlighted several shortcomings of traditional tokenization methods, including low vocabulary utilization. In response, alternative approaches like FSQ have been explored (Mentzer et al., 2024, Yu et al., 2023b). These discrete visual tokens facilitate integration with next-token prediction in language models and multimodal models (Yu et al., 2023b, Liu et al., 2024a, Xie et al., 2024, Wang et al., 2024). Parallel to this, other studies have investigated the use of continuous embeddings for representing images and videos (Lin et al., 2023, Liu et al., 2023, Zhu et al., 2023). Other research proposes next-scale prediction (Tian et al., 2024), which employs hierarchical multi-scale modeling to first capture coarse details, followed by finer ones. However, this approach requires a manually defined hierarchy and uses a fixed number of tokens. Our work, while complementary to these efforts, focuses on adaptive representation and can be directly applied to both discrete and continuous approaches. Refer to caption Figure 9: Comparison of token usage versus frequency magnitude in single-block and multi-block frequency analysis (ElasticTok-FSQ). Both scatter plots show a strong positive correlation between frequency magnitude and token usage in a single-block setting a multi-block setting. The red lines represent the linear regression fits for each case. Compression. Adaptive learning for images and video have also been well-studied in the context of variable-rate compression. JPEG (Christopoulos et al., 2000) remains one of the most popular lossy image compression algorithms in the world, using a combination of DCT with quantization followed by entropy coding. For video, codecs such as H264 (Wiegand et al., 2003), H265 (Sze et al., 2014), VP9 (Mukherjee et al., 2015), and AV1 (Han et al., 2021) are popularly used compression algorithms. Prior works have additionally explored extending neural networks to learn more effective compressors. Theis et al. (2022) and Minnen et al. (2018) introduce methods that leverage CNNs to learn effective variable bit-rate compression algorithms over images competitive with JPEG. Li et al. (2023a), Mentzer et al. (2022), and Li et al. (2021) extend neural networks to learn how to effectively compress videos. While similar to our work in that these methods also learn adaptive encodings, these compression methods are more difficult to utilize for downstream training and generation tasks due to lack of direct compatibility. In contrast, our work builds upon existing tokenization strategies (VAE, FSQ) that have shown to work well for such downstream tasks, as well as take advantage of the adaptive representations to better utilize compute.
@@ 6 Discussion and Conclusion
In this work, we propose an elastic representation approach to address the inefficiencies of traditional video encoding approaches through an adaptive encoding method that selectively encodes new information based on the context of previous frames. By dynamically allocating resources, it reduces computational costs while maintaining high-quality video representation. The proposed technique of dropping tokens at the end of each sequence allows the model to prioritize essential information, ensuring scalability and efficiency during inference. ElasticTok demonstrates strong performance in both video representation and downstream tasks. Lastly, we identify some limitations of our method, as well as possible directions for future work. Masking schemes: Our model currently slightly underperforms baselines at the tail ends of encoding lengths, which we hypothesize may be partially due to conflicting representations that the model needs to learn in each case (low frequency, global encodings vs high frequency, local encodings). Preliminary investigations showed that changing the way tokens are distributed (as opposed to our method of left-aligned tokens) improved reconstruction accuracy. Future work may also explore learnable encoding masks to dynamically select which tokens to keep for each input. Other temporal modalities: Although our primary focus was on image and video, our method is generally modality-agnostic and can be extended to any temporal data, such as audio and decision-making trajectories (e.g., state, action, reward). Conditional encoding and different training objectives: Our focus was on leveraging temporal encoding to achieve more efficient tokenization while retaining high reconstructed visual quality. However, our method can be explored more broadly as a means of encouraging meaningful lossy encodings. For example, robotics models can learn text-conditioned visual encoders that capture only task-relevant information from an image. In other scenarios, different reconstruction objectives could prioritize facial or textual reconstruction while ignoring background reconstruction accuracy.
@@ Acknowledgments
We gratefully acknowledge the Google TPU Research Cloud (TRC) program for providing access to TPUs. Pieter Abbeel holds concurrent appointments as a Professor at UC Berkeley and as an Amazon Scholar. This paper describes work performed at UC Berkeley and is not associated with Amazon. We extend our thanks to Xinyang Geng, Jimmy Shi, Tim Rocktäschel, Joao Carreira, and Satinder Singh for their valuable discussions and insightful feedback on an earlier version of this paper, and to Doina Precup for valuable support of this work.
@@ Appendix A Model configuration
ElasticTok-VAE (Continuous) ElasticTok-FSQ (Discrete) Parameters 210M 210M Frame Resolution $256\times 256$ $256\times 256$ Block Size ($M_{max}$) 2048 tokens (4 frames) 4096 tokens (4 frames) $M_{min}$ 128 tokens 256 tokens Max Number of Frames 1024 1024 Patch Size (T, H, W) (2, 8, 8) (1, 8, 8) Hidden Size 1024 1024 FFN Size 2048 2048 Encoder Layers 10 10 Decoder Layers 10 10 RoPE Theta 5000000 50000000 Max Sequence Length 512K 1M FSQ Dims N/A 8, 8, 8, 5, 5, 5 (64k codes) VAE Dim 8 N/A KL Weight 1e-8 N/A